OpenStudy (anonymous):

Multiply. Assume that all expressions are defined. \[\frac{ x ^{2}-16 }{ x ^{2}-4x+4 } \times \frac{ x-2 }{ x ^{2}+6x+8 }\]

OpenStudy (anonymous):

ok x^2 -16 can be simplified into (x-4)(x+4) using (a+b)(a-b) = a^2 -b^2 x^2-4x+4 = (x-2)^2 using (a-b)^2 = a^2 -2ab+b^2 x^2+6x+8 = x^2 + 4x +2x + 8 = x(x+4)+2(x+4) = (x+2)(x+4) now we have so far : (x-4)(x+4) * (x-2) ---------- ------ (x-2)^2 (x+2)(x+4)

OpenStudy (anonymous):

can you do it now ?

OpenStudy (anonymous):

:/ no

OpenStudy (anonymous):

ok you see we have x+4 in the denominator and in the numerator so we can eliminate smae with (x-2) and (x-2)^2 we will left with x-2 in the denominator

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-4)/(x-2)(x+2)

OpenStudy (anonymous):

so its x-4/x+2

OpenStudy (anonymous):

no i wrote you the final answer in the drawing

OpenStudy (anonymous):

there is (x-2) in the denominator since it was squared and on the numerator it was to the power of one so there is one left in the denominator..

OpenStudy (anonymous):

okay thank you!!!

OpenStudy (anonymous):

if a = x-2 so its like : \[\frac{ a }{ a^2 } = \frac{ 1 }{ a }\]

OpenStudy (anonymous):

yw